# Centered Moments of Weighted One-Level Densities of $GL(2)$ $L$-Functions

### v3

@@ Abstract.
This study took place at the SMALL REU at Williams College and was supported by Williams College and the National Science Foundation (Grant DMS2241623). The authors are also grateful for the support of Columbia University, Princeton University, University of Michigan, University of Washington, University of Wisconsin, Williams College, and Yale University. The authors are also grateful to Andrew Knightly for suggesting improvements to the paper.
@@ 1. Introduction
In 1972, Montgomery [Mon73] initiated the connection between number theory and random matrix theory, proving that, for suitable test functions, the pair correlation of the zeros of the Riemann zeta function $\zeta(s)$ agrees with the pair correlation of the eigenvalues of the Gaussian Unitary Ensemble (GUE). In his seminal work, he furthermore conjectured a correspondence between any local statistic of the zeros of $\zeta(s)$ and the eigenvalues of the GUE. In 1987, Odlyzko [Odl87, Odl01] provided numerical evidence for this conjecture, verifying the correspondence for particular local statistics: the pair correlation and the nearest neighbor spacing distribution. Subsequent works by Hejhal [Hej94] on the triple correlation of $\zeta(s)$ and Rudnick and Sarnak [RS96] on the $n$-level correlation of any automorphic cuspidal $L$-function, all for suitably restricted test functions, provided strong evidence for Montgomery’s conjecture in generality. These studies suggested a remarkable universality in number theory: the statistical profile of any $L$-function matches that of one and only one of the numerous random matrix ensembles, the GUE.
@@
At the time, however, there were a few reasons to believe that this conjecture did not capture the connection in full. For one, certain local statistics, such as the $n$-level correlation, are insensitive to the behavior of any finite set of zeros. As there are many important problems in number theory concerning only finite sets of zeros (e.g., the Birch and Swinnerton-Dyer conjecture [BS63, BS65] is concerned only with the low-lying zeros of elliptic curve $L$-functions), this marked a considerable shortcoming of the conjecture on the number theory side. Katz and Sarnak [KS99a, KS99] confirmed that indeed more attention is needed in this regard. They showed that the $n$-level correlation of eigenvalues is the same coming from the GUE and all five classical compact groups (unitary, symplectic, and orthogonal, split and unsplit by sign). They addressed this possibility of confounding by defining a new statistic called the one-level density, which is not only distinguishable across the classical compact groups but also sensitive to changes in low-lying zeros (near the central point). For any $L$-function $L(s,f)$, denoting its non-trivial zeros by $\rho_{f}=1/2+i\gamma_{f}$ [^1] If the Grand Riemann Hypothesis (GRH) is true, then $\gamma$ would be real, and the non-trivial zeros $\rho$ could be ordered accordingly. There are interesting interpretations of this ordering in the context of random matrix theory and nuclear physics, actuating connections to the eigenvalues of classical compact groups and the energy levels of heavy nuclei. With this being said, we do not assume the truth of the GRH in what follows, performing our study in generality., they defined its one-level density by
$D(f,\phi)\ :=\ \sum_{\rho_{f}}\phi\left(\frac{\gamma_{f}\log Q_{f}}{2\pi}\right),$ (1.1)
@@
where $\phi$ is an even Schwartz (test) function, and $Q_{f}$ is the analytic conductor of $f$. The low-lying zeros of $L(s,f)$ have imaginary part approximately $1/\log Q_{f}$ and the average spacing between zeros with imaginary part $T$ is known to be approximately $1/\log T$. Hence, the scaled zeros, $\frac{\gamma_{f}\log Q_{f}}{2\pi}$, have average spacing $1$, meaning there are only finitely many up to any given constant. As it is not instructive to statistically survey the finitely many low-lying zeros of a single $L$-function, Katz and Sarnak passed to the study of collections of naturally related (similarly behaved) $L$-functions, “families” so to speak. In doing so, they were able to execute averages and take limits, as is customary in analytic number theory, to identify the common behavior underlying the $L$-functions in the family.
@@
$\mathcal{A}^{w}_{\mathcal{C}}(Q)\ :=\ \frac{\sum_{c\in\mathcal{C}}w(c)Q(c)}{\sum_{c\in\mathcal{C}}w(c)}.$ (1.2)
For unweighted averages, i.e., when $w$ is identically $1$ on $\mathcal{C}$, we omit the superscript in the notation (1.2).
@@
$\mathcal{A}_{\mathcal{F}_{k}}(D(\cdot,\phi))\ :=\ \frac{\sum_{L(s,f)\in\mathcal{F}_{k}}D(f,\phi)}{|\mathcal{F}_{n}|}.$ (1.3)
@@
$\lim_{k\to\infty}\mathcal{A}_{\mathcal{F}_{k}}(D(\cdot,\phi))\ =\ \int_{-\infty}^{\infty}\phi(x)W_{G}(x)dx.$ (1.4)
@@
$W_{\text{U}}(x)\$= $=\ 1,$ (1.5)
$W_{\text{Sp}}(x)\$= $=\ 1-\frac{\sin(2\pi x)}{2\pi x},$ (1.6)
$W_{\text{SO(even)}}(x)\$= $=\ 1+\frac{\sin(2\pi x)}{2\pi x},$ (1.7)
$W_{\text{SO(odd)}}(x)\$$=\ \delta_{0}+1-\frac{\sin(2\pi x)}{2\pi x},$ (1.8)
$W_{\text{O}}(x)\$= $=\ 1+\frac{1}{2}\delta_{0}(x),$ (1.9)
@@
$D_{n}(f;\phi_{1},\dots,\phi_{n})\ :=\ \sum_{\begin{subarray}{c}\rho_{f}^{(1)},\dots,\rho_{f}^{(n)}\\ \gamma_{f}^{(i)}\neq\pm\gamma_{f}^{(j)}\end{subarray}}\phi_{1}\left(\frac{\gamma_{f}^{(1)}\log Q_{f}}{2\pi}\right)\cdots\phi_{n}\left(\frac{\gamma_{f}^{(n)}\log Q_{f}}{2\pi}\right).$ (1.10)
The $n$-level density of a family is usually calculated in terms of its one-level density using the principle of inclusion-exclusion [Rub01, Gao08]. This approach, however, relies on our knowledge of the distribution of the signs of the functional equations in the family, which is beyond current theory. In this view, Hughes and Rudnick [HR02] initiated the study of a more tractable $n$-level statistic: the $n$th centered moment of the one-level density. More precisely, for the general sub-family $\mathcal{F}_{k}$ of $L$-functions given above, the $n$th centered moment of its one-level density is defined as
$\mathcal{A}_{\mathcal{F}_{k}}\left[\left(D(\cdot,\phi)-\mathcal{A}_{\mathcal{F}_{k}}(D(\cdot,\phi))\right)^{n}\right]\ :=\ \frac{\sum_{L(s,f)\in\mathcal{F}_{k}}\left(D(f,\phi)-\mathcal{A}_{\mathcal{F}_{k}}(D(\cdot,\phi))\right)^{n}}{|\mathcal{F}_{k}|}.$ (1.11)
@@
The general approach in calculating these density statistics for families of $L$-functions is to find first an explicit formula converting the sum over zeros in (1.1) to a sum over primes, then an asymptotic (trace) formula for the resulting sum over primes. Finding the explicit formula is usually straightforward, e.g., [ILS00, Lemma 4.1], but finding the trace formula usually entails intricate equidistribution laws governing the family. In the specific case of families of automorphic forms, trace formulae naturally involve weights. For example, harmonic weights naturally arise in the Petersson trace formula for families of cuspidal newforms [ILS00, (2.53)], and analytic and harmonic weights naturally arise in the Kuznetsov and Selberg trace formulae for families of Maass forms [KL13, GK12, Remark 1.6]. In this view, many statistical surveys on families of automorphic $L$-functions [Alp+15, GK12, ILS00] maintain these weights and calculate weighted, rather than unweighted, densities of their zeros. In all the above surveys, the weights have proved to be innocuous, in that they do not affect the symmetry type of the low-lying zeros, i.e., the uniform and weighted zeros have the same limiting distribution. However, this is not true in general.
@@
  • $\mathcal{F}_{1}=\bigcup\limits_{k}\mathcal{F}_{k}(1)^{\prime}$ with $k$ ranging over even integers satisfying $\tau(\chi)^{2}\neq-i^{k}D$;
  • $\mathcal{F}_{2}=\bigcup\limits_{k,N}\mathcal{F}_{k}(N)^{\prime}$ with $k>2$ ranging over even integers satisfying $\tau(\chi)^{2}=-i^{k}D$, and $N$ ranging over primes not dividing $rD$.
@@
$w_{\chi,r}(f)\ :=\ \frac{\Lambda\left(\frac{1}{2},f\times\chi\right)|a_{f}(r)|^{2}}{\lVert f\rVert^{2}},$ (1.12)
@@
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\ :=\ \frac{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)D(f,\phi)}{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)}.$ (1.13)
They proved that for suitably restricted test functions, the limiting distribution of the average weighted one-level densities for both $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ varies depending on the triviality of the twisting character in the weight (1.12):
$\lim_{k\to\infty}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(1)^{\prime}}\left(D(\cdot,\phi_{1})\right)\$$=\ \lim_{k+N\to\infty}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(D(\cdot,\phi_{1})\right)$
$=\ \begin{cases}\int_{-\infty}^{\infty}\phi_{1}(x)W_{\text{Sp}}(x)dx&\text{if }\chi\text{ trivial},\\ \int_{-\infty}^{\infty}\phi_{1}(x)W_{\text{O}}(x)dx&\text{if }\chi\text{ non-trivial};\end{cases}$ (1.14)
here, the limits are over the aforementioned values of $k$ and $N$. It is instructive to compare (1) to [ILS00, Theorem 1.1], which establishes that the limiting distribution of the average unweighted one-level densities for similar families of Hecke cuspidal newforms is orthogonal. In the case of the trivial twisting character, the twisted $L$-function is equal to the original $L$-function; hence the central value of the (twisted) $L$-function is strongly correlated to the zeros near the central point of the original $L$-function (a large central $L$-value generally implies few low-lying zeros). This explains why trivial weighting characters change the limiting distribution from orthogonal to symplectic. On the other hand, in the case of non-trivial twisting characters, the twisted $L$-function (in particular, the twisted central value) is presumably independent of the original $L$-function (in particular, its low-lying zeros). This is to say that these non-trivial characters make for somewhat random weights, independent of the low-lying zeros that they are weighting, so the distribution, in the limit, converges as expected to the orthogonal distribution.
@@
$:=\ \frac{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)\left(D(f,\phi)-\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\right)^{n}}{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)}.$ (1.15)
In addition to the notation so far, we adopt that for any test function $\phi$,
$\sigma_{\phi}\ :=\ \left(\int_{-\infty}^{\infty}\widehat{\phi}^{2}(y)|y|\,dy\right)^{1/2}.$ (1.16)
Our main result is that for suitably restricted test functions $\phi$, the weighted $n$th centered moment of the average one-level densities for both $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ converge to the $n$th centered moment of a Gaussian distribution with variance $\sigma_{\phi}^{2}$.
@@
Let $\chi$ be a real, primitive Dirichlet character of modulus $D\geq 1$, and let $r$ be a positive integer relatively prime to $D$. For any positive integer $n$ and any test functions $\phi$ with $\text{supp}(\widehat{\phi})\subset\left(-\frac{1}{2n},\frac{1}{2n}\right)$,
$\lim_{k\to\infty}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(1)^{\prime}}\left[\left(D(\cdot,\phi)-\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(1)^{\prime}}(D(\cdot,\phi))\right)^{n}\right]\ =\$$\begin{cases}(n-1)!!\ \sigma_{\phi}^{n}&\text{if }n\text{ even},\\ 0&\text{if }n\text{ odd},\end{cases}$ (1.17)
$\lim_{k+N\to\infty}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(D(\cdot,\phi)-\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\right)^{n}\right]\ =\$$\begin{cases}(n-1)!!\ \sigma_{\phi}^{n}&\text{if }n\text{ even},\\ 0&\text{if }n\text{ odd}.\end{cases}$ (1.18)
@@
Note that we recover the result of Knightly and Reno in the case $n=1$. The right-hand side would equal 0 in this case, and moving over the average weighted one-level density from the left-hand side returns us to (1).
Remark 1.3.
It would be interesting to increase the support beyond $(-\frac{1}{n},\frac{1}{n})$ because we can start to differentiate $W_{\text{O}}(x)$ and $W_{\text{Sp}}(x)$ as distributions only beyond this range, as noted in the opening discussion. In the current scope, Theorem 1.1 does not detect the difference between the orthogonal and symplectic distributions, and hence the dependence between weights (particularly the triviality of the twisting character therein) and symmetry type.
@@
When $n=1$, we recover from Theorem 1.1 one of the two main results from the study of Knightly and Reno: (1). In their other main result [KR19, Theorem 1.3], Knightly and Reno considered different weights:
$w_{\chi}^{\prime}(f)=\frac{\Lambda\left(\frac{1}{2},f\times\chi\right)\Lambda\left(\frac{1}{2},f\right)}{\lVert f\rVert^{2}}.$ (1.19)
@@
In the case of trivial twisting characters, it is not obvious a priori that the number theoretic terms resulting from our analysis match up exactly with the expected random matrix theoretic terms (1.18). This apparent mismatch is observed in most, if not all, $n$-level calculations [Rub01, Gao08, HM07, HR02, Sos00]. This is because the random matrix theoretic terms are derived for arbitrary support, cf. [KS99, KS99a]; whereas the number theoretic terms are derived for restricted support, out of technical necessity. Besides, since these calculations usually use some combinatorial argument (such as the principle of inclusion-exclusion) to express the $n$-level statistic in terms of the corresponding one-level statistic, it is instructive to reverse this combinatorial argument in the end, i.e., unravel the combinatorial expression that emerges after the one-level statistic is calculated. In this vein, Soshnikov [Sos00] developed a combinatorial trick relating to generating series, which has been referenced and adapted in several subsequent $n$-level calculations [HM07, HR02]; the essence of this trick is a deep combinatorial fact called the Hunt-Dyson formula. Novelly, our combinatorial strategy for unraveling the number theoretic terms in Subsection 5.2 does not rely on this trick. We use only elementary methods to rewrite the resulting combinatorial factors and motivate our otherwise unassuming applications of the binomial theorem. This represents the main challenge in going from Knightly and Reno’s one-level calculation, which did not require any combinatorial argument, to our $n$-level calculation. It would be interesting to assess the applicability of our elementary combinatorial argument in other $n$-level calculations, like the ones referenced above.
@@
In view of the many incentives to extending support, e.g., Remarks 1.3 and 1.4, it is worth mentioning that the current support restrictions are in place only to bound the contribution of the error term in the asymptotic trace formula (see Lemma 5.1). All other contributions are analyzed unconditionally, for arbitrary support. Extending the support, therefore, entails finding more lower-order terms in the respective trace formula and thereby analyzing a smaller error term.

### v4

@@ Abstract.
This study took place at the SMALL REU at Williams College and was supported by Williams College and the National Science Foundation (Grant DMS2241623). The authors are also grateful for the support of Columbia University, Princeton University, University of Michigan, University of Washington, University of Wisconsin, Williams College, and Yale University. We thank Andrew Knightly for suggesting improvements to the paper, and also our journal referee for detailed and helpful comments.
@@ 1. Introduction
In 1972, Montgomery [Mon73] formalized the connection between number theory and random matrix theory, proving that, for suitable test functions, the pair correlation of the zeros of the Riemann zeta function $\zeta(s)$ agrees with the pair correlation of the eigenvalues of the Gaussian Unitary Ensemble (GUE). In his seminal work, he furthermore conjectured a correspondence between any local statistic of the zeros of $\zeta(s)$ and the eigenvalues of the GUE. In 1987, Odlyzko [Odl87, Odl01] provided numerical evidence for this conjecture, verifying the correspondence for particular local statistics: the pair correlation and the nearest neighbor spacing distribution. Subsequent works by Hejhal [Hej94] on the triple correlation of $\zeta(s)$ and Rudnick and Sarnak [RS96] on the $n$-level correlation of any automorphic cuspidal $L$-function, all for suitably restricted test functions, provided strong evidence for Montgomery’s conjecture in generality. These studies suggested a remarkable universality in number theory: the statistical profile of any $L$-function matches that of one and only one of the numerous random matrix ensembles, the GUE.
@@
At the time, however, there were a few reasons to believe that this conjecture did not capture the connection in full. For one, certain local statistics, such as the $n$-level correlation, are insensitive to the behavior of any finite set of zeros. As there are many important problems in number theory concerning only finite sets of zeros (e.g., the Birch and Swinnerton-Dyer conjecture [BS63, BS65] is concerned only with the low-lying zeros of elliptic curve $L$-functions), this marked a considerable shortcoming of the conjecture on the number theory side. Katz and Sarnak [KS99, KS99a] confirmed that indeed more attention is needed in this regard. They showed that the $n$-level correlation of eigenvalues is the same coming from the GUE and all five classical compact groups (unitary, symplectic, and orthogonal, split and unsplit by sign). They addressed this possibility of confounding by defining a new statistic called the one-level density, which is not only distinguishable across the classical compact groups but also sensitive to changes in low-lying zeros (near the central point). For any $L$-function $L(s,f)$, denoting its non-trivial zeros by $\rho_{f}=1/2+i\gamma_{f}$ [^1] If the Grand Riemann Hypothesis (GRH) is true, then $\gamma$ would be real, and the non-trivial zeros $\rho$ could be ordered accordingly. There are interesting interpretations of this ordering in the context of random matrix theory and nuclear physics, actuating connections to the eigenvalues of classical compact groups and the energy levels of heavy nuclei. With this being said, we do not assume the truth of the GRH in what follows, performing our study in generality., they defined its one-level density by
(1.1) $D(f,\phi)\ :=\ \sum_{\rho_{f}}\phi\left(\frac{\gamma_{f}\log Q_{f}}{2\pi}\right),$
@@
where $\phi$ is an even Schwartz (test) function, and $Q_{f}$ is the analytic conductor of $f$. The low-lying zeros of $L(s,f)$ have imaginary part approximately $1/\log Q_{f}$ and the average spacing between zeros with imaginary part $T$ is known to be approximately $1/\log T$. Hence, the scaled zeros, $\frac{\gamma_{f}\log Q_{f}}{2\pi}$, have average spacing $1$, meaning there are only finitely many up to any real number $C>0$. As it is not instructive to statistically survey the finitely many low-lying zeros of a single $L$-function, Katz and Sarnak passed to the study of collections of naturally related (similarly behaved) $L$-functions— “families” so to speak. In doing so, they were able to execute averages and take limits, as is customary in analytic number theory, to identify the common behavior underlying the $L$-functions in the family.
@@
(1.2) $\mathcal{A}^{w}_{\mathcal{C}}(Q)\ :=\ \frac{\sum_{c\in\mathcal{C}}w(c)Q(c)}{\sum_{c\in\mathcal{C}}w(c)}.$
For unweighted averages, i.e., when $w$ is identically $1$ on $\mathcal{C}$, we omit the superscript in (1.2).
@@
(1.3) $\mathcal{A}_{\mathcal{F}_{k}}(D(\cdot,\phi))\ :=\ \frac{\sum_{L(s,f)\in\mathcal{F}_{k}}D(f,\phi)}{|\mathcal{F}_{k}|}.$
@@
(1.4) $\lim_{k\to\infty}\mathcal{A}_{\mathcal{F}_{k}}(D(\cdot,\phi))\ =\ \int_{-\infty}^{\infty}\phi(x)W_{G}(x)dx.$
@@
(1.5) $W_{\text{U}}(x)\$= $=\ 1,$
(1.6) $W_{\text{Sp}}(x)\$= $=\ 1-\frac{\sin(2\pi x)}{2\pi x},$
(1.7) $W_{\text{SO(even)}}(x)\$= $=\ 1+\frac{\sin(2\pi x)}{2\pi x},$
(1.8) $W_{\text{SO(odd)}}(x)\$$=\ \delta_{0}+1-\frac{\sin(2\pi x)}{2\pi x},$
(1.9) $W_{\text{O}}(x)\$= $=\ 1+\frac{1}{2}\delta_{0}(x),$
@@
(1.10) $D_{n}(f;\phi_{1},\dots,\phi_{n})\ :=\ \sum_{\begin{subarray}{c}\rho_{f}^{(1)},\dots,\rho_{f}^{(n)}\\ \gamma_{f}^{(i)}\neq\pm\gamma_{f}^{(j)}\end{subarray}}\phi_{1}\left(\frac{\gamma_{f}^{(1)}\log Q_{f}}{2\pi}\right)\cdots\phi_{n}\left(\frac{\gamma_{f}^{(n)}\log Q_{f}}{2\pi}\right).$
The $n$-level density of a family is usually calculated in terms of its one-level density using the principle of inclusion-exclusion [Rub01, Gao13]. This approach, however, relies on our knowledge of the distribution of the signs of the functional equations in the family, which is not known in general. [^2] For the family of weight $k$, level $N$ cuspidal newforms that we study in this paper, the distribution of the signs of the functional equations has been formulated by Martin [Mar23, Theorem 1.2]. Using this, it would be interesting to consider extending our results here from $n$th centered moments of the one-level density to $n$-level densities. In this view, Hughes and Rudnick [HR02] initiated the study of a more tractable $n$-level statistic: the $n$th centered moment of the one-level density. More precisely, for the general sub-family $\mathcal{F}_{k}$ of $L$-functions given above, the $n$th centered moment of its one-level density is defined as
(1.11) $\mathcal{A}_{\mathcal{F}_{k}}\left[\left(D(\cdot,\phi)-\mathcal{A}_{\mathcal{F}_{k}}(D(\cdot,\phi))\right)^{n}\right]\ :=\ \frac{\sum_{L(s,f)\in\mathcal{F}_{k}}\left(D(f,\phi)-\mathcal{A}_{\mathcal{F}_{k}}(D(\cdot,\phi))\right)^{n}}{|\mathcal{F}_{k}|}.$
@@
The general approach in calculating these density statistics for families of $L$-functions is to first find an explicit formula converting the sum over zeros in (1.1) to a sum over primes, then an asymptotic (trace) formula for the resulting sum over primes. Finding the explicit formula is usually straightforward, e.g., [ILS00, Lemma 4.1], but finding the trace formula usually entails intricate equidistribution laws governing the family. In the specific case of families of automorphic forms, trace formulae naturally involve weights. For example, harmonic weights naturally arise in the Petersson trace formula for families of cuspidal newforms [ILS00, (2.53)], and analytic and harmonic weights naturally arise in the Kuznetsov and Selberg trace formulae for families of Maass forms [KL13, GK12, Remark 1.6]. In this view, many statistical surveys on families of automorphic $L$-functions [Alp+15, GK12, ILS00] maintain these weights and calculate weighted, rather than unweighted, densities of their zeros. In all the above surveys, the weights have proved to be innocuous, in that they do not affect the symmetry type of the low-lying zeros, i.e., the uniform and weighted zeros have the same limiting distribution. However, this is not true in general.
@@
  • $\mathcal{F}_{1}=\bigcup_{k}\mathcal{F}_{k}(1)^{\prime}$ with $k$ ranging over even integers satisfying $\tau(\chi)^{2}\neq-i^{k}D$;
  • $\mathcal{F}_{2}=\bigcup_{k,N}\mathcal{F}_{k}(N)^{\prime}$ with $k>2$ ranging over even integers satisfying $\tau(\chi)^{2}=-i^{k}D$, and $N$ ranging over primes not dividing $rD$. We order the sub-families $\mathcal{F}_{k}(N)^{\prime}$ so that $k+N$ is increasing.
@@
(1.12) $w_{\chi,r}(f)\ :=\ \frac{\Lambda\left(\frac{1}{2},f\times\chi\right)|a_{f}(r)|^{2}}{\lVert f\rVert^{2}},$
@@
(1.13) $\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\ :=\ \frac{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)D(f,\phi)}{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)}.$
They proved that for suitably restricted test functions $\phi$, the limiting distribution of the average weighted one-level densities for both $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ varies depending on the triviality of the twisting character in the weight (1.12):
$\lim_{k\to\infty}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(1)^{\prime}}\left(D(\cdot,\phi)\right)\$$=\ \lim_{k+N\to\infty}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left(D(\cdot,\phi)\right)$
(1.14) $=\ \begin{cases}\int_{-\infty}^{\infty}\phi(x)W_{\text{Sp}}(x)dx&\text{if }\chi\text{ trivial},\\ \int_{-\infty}^{\infty}\phi(x)W_{\text{O}}(x)dx&\text{if }\chi\text{ non-trivial},\end{cases}$
where the limits are over the aforementioned values of $k$ and $N$. It is instructive to compare (1) to [ILS00, Theorem 1.1], which establishes that the limiting distribution of the average unweighted one-level densities for similar families of Hecke cuspidal newforms is orthogonal. In the case of the trivial twisting character, the twisted $L$-function is equal to the original $L$-function; hence the central value of the (twisted) $L$-function is strongly correlated to the zeros near the central point of the original $L$-function (a large central $L$-value generally implies few low-lying zeros). This explains why trivial weighting characters change the limiting distribution from orthogonal to symplectic. On the other hand, in the case of non-trivial twisting characters, the twisted $L$-function (in particular, the twisted central value) is presumably independent of the original $L$-function (in particular, its low-lying zeros). This is to say that these non-trivial characters make for somewhat random weights, independent of the low-lying zeros that they are weighting, so the distribution, in the limit, converges as expected to the orthogonal distribution.
@@
(1.15) $:=\ \frac{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)\left(D(f,\phi)-\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\right)^{n}}{\sum_{f\in\mathcal{F}_{k}(N)^{\prime}}w_{\chi,r}(f)}.$
In addition to the notation so far, we define for any test function $\phi$,
(1.16) $\sigma_{\phi}\ :=\ \left(\int_{-\infty}^{\infty}\widehat{\phi}(y)^{2}|y|\,dy\right)^{1/2}.$
Our main result is that for suitably restricted test functions $\phi$, the weighted $n$th centered moment of the average one-level densities for both $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ converges to the $n$th centered moment of a Gaussian distribution with variance $\sigma_{\phi}^{2}$.
@@
Let $\chi$ be a real, primitive Dirichlet character of modulus $D\geq 1$, and let $r$ be a positive integer relatively prime to $D$. For any positive integer $n$ and any test function $\phi$ with $\text{supp}(\widehat{\phi})\subset\left(-\frac{1}{2n},\frac{1}{2n}\right)$,
(1.17) $\lim_{k\to\infty}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(1)^{\prime}}\left[\left(D(\cdot,\phi)-\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(1)^{\prime}}(D(\cdot,\phi))\right)^{n}\right]\ =\$$\begin{cases}(n-1)!!\ \sigma_{\phi}^{n}&\text{if }n\text{ even},\\ 0&\text{if }n\text{ odd},\end{cases}$
(1.18) $\lim_{k+N\to\infty}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(D(\cdot,\phi)-\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}(D(\cdot,\phi))\right)^{n}\right]\ =\$$\begin{cases}(n-1)!!\ \sigma_{\phi}^{n}&\text{if }n\text{ even},\\ 0&\text{if }n\text{ odd}.\end{cases}$
@@
It would be interesting to increase the support beyond $(-\frac{1}{n},\frac{1}{n})$ because we can start to distinguish between $W_{\text{O}}(x)$ and $W_{\text{Sp}}(x)$ as distributions only beyond this range, as noted in the opening discussion. In the current scope, Theorem 1.1 does not detect the difference between the orthogonal and symplectic distributions, and hence the dependence between weights (particularly the triviality of the twisting character therein) and symmetry type.
@@
When $n=1$, we use one of the two main results from the study of Knightly and Reno: (1). In their other main result [KR19, Theorem 1.3], Knightly and Reno considered different weights:
(1.19) $w_{\chi}^{\prime}(f)=\frac{\Lambda\left(\frac{1}{2},f\times\chi\right)\Lambda\left(\frac{1}{2},f\right)}{\lVert f\rVert^{2}}.$
@@
In the case of trivial twisting characters, it is not obvious a priori that the number theoretic terms resulting from our analysis match up exactly with the expected random matrix theoretic terms (1.18). This apparent mismatch is observed in most, if not all, $n$-level calculations [Rub01, Gao13, HM07, HR02, Sos00]. This is because the random matrix theoretic terms are derived for arbitrary support, cf. [KS99, KS99a]; whereas the number theoretic terms are derived for restricted support, out of technical necessity. Besides, since these calculations usually use some combinatorial argument (such as the principle of inclusion-exclusion) to express the $n$-level statistic in terms of the corresponding one-level statistic, it is instructive to reverse this combinatorial argument in the end, i.e., unravel the combinatorial expression that emerges after the one-level statistic is calculated. In this vein, Soshnikov [Sos00] developed a combinatorial trick relating to generating series, which has been referenced and adapted in several subsequent $n$-level calculations [HM07, HR02]; the essence of this trick is a deep combinatorial fact called the Hunt-Dyson formula. Novelly, our combinatorial strategy for unraveling the number theoretic terms in Subsection 5.2 does not rely on this trick. We use only elementary methods to rewrite the resulting combinatorial factors and motivate our otherwise unassuming applications of the binomial theorem. This represents the main challenge in going from Knightly and Reno’s one-level calculation, which did not require any combinatorial argument, to our $n$-level calculation. It would be interesting to assess the applicability of our elementary combinatorial argument in other $n$-level calculations, like the ones referenced above.
@@
In view of the many incentives to extending support, e.g., Remarks 1.2 and 1.3, it is worth mentioning that the current support restrictions are in place only to bound the contribution of the error term in the asymptotic trace formula (see Lemma 5.1). All other contributions are analyzed unconditionally, for arbitrary support. Extending the support, therefore, entails finding more lower-order terms in the respective trace formula and thereby analyzing a smaller error term.